Find the monthly payment on an auto loan of 200000 to be akortized over a 5-year period at a rate of 9%.
1
Expert's answer
2021-08-30T16:23:02-0400

Given Information :

Auto loan - 200000

Period - 5 years

Rate - 9%

Calculation of the monthly payment on an auto loan:

Present value = Monthly amount × "[(1−\frac{\frac{1}{(1+r)^

n}}{

r}]"

where, r = 9%/12  = 0.75%

n = 5 ×12 = 60 months

So, the amount will be :

"200000 = Monthly \space amount ×[(1−\frac{\frac{1}{(1+0.0075)^{

60}}}{0.0075}]"

200000 = Monthly amount ×48.17337352093

Monthly amount = 200000/48.17337352093

Monthly amount = 4151.67 or 4152


Therefore, the monthly payment on a auto loan is 4151.67 or 4152